OpenStudy (anonymous):

The manager of a sandwich shop is planning to make a circle graph showing the types of sandwiches sold in one day. The table below summarizes these data. Sandwiches Sold in One Day Roast beef 195 Turkey-and-cheese 125 Ham-and-cheese 75 Vegetable 50 Other 55 What central angle should the manager use for the section representing the ham-and-cheese sandwich? A 270° B 75° C 54° D 15°

Directrix (directrix):

C) 54° ------- 500 sandwiches were sold. Of these, 75 were ham and cheese. Therefore, 75 is to 500 as x is to 360 where x is the central angle of the circle graph 75/500 = x/360 x = 54 degrees